### v1

@@ 6.1.1 Optimising survey strategy by increasing sky area
In this simple test we let the parameters $\bm{\theta}\,{=}\,\{b_{\textrm{H{i}}},f,\alpha_{\perp},\alpha_{\parallel},f_{\text{NL}}\}$ vary, whilst assuming everything else fixed. We extract $H(z)$ from the AP parameter $\alpha_{\parallel}(z)$ (see Equation 26), and marginalise over $b_{\textrm{H{i}}}$ and $\alpha_{\perp}$. The parameters $b_{\textrm{H{i}}}$ and $f$ should be fully degenerate in linear theory for the monopole of the Hi power spectrum. For intensity mapping, it has been shown that going to higher order multipoles is especially crucial for breaking such degeneracies [28, 86, 114], so in this analysis we include the quadrupole and hexadecapole, i.e. $\bm{P}_{\ell}\,{=}\,\{P_{0},P_{2},P_{4}\}$, which assumes observational anisotropies can be well controlled [32, 54, 130]. There will be additional degeneracies with $\overline{T}_{\textrm{H{i}}}$ and $\sigma_{8}$, for which we fix here, thus we are inherently assuming they have been independently well measured, or alternatively constraints would include these factors as degenerate quantities, e.g. $f\sigma_{8}\overline{T}_{\textrm{H{i}}}$. The $\sigma_{8}$ degeneracy is a general problem for large-scale structure, but the $\overline{T}_{\textrm{H{i}}}$ is, of course, more unique to intensity mapping [identified in 28, 43]. Exploiting tomography, i.e. levering the differing redshift evolution of degenerate parameters across bins, can provide some immediate relief here. Additionally, 1-point statistics, such as the voxel intensity distribution (VID) [34, 120, 27], can capture the non-Gaussian information in the line intensity maps, thus combining the VID with the power spectrum helps to break degeneracies between astrophysical and cosmological parameters [115]. Furthermore, going to non-linear scales with interferometers, where scale dependence will enter and since Hi clustering should be accurately described by perturbative methods at these mildly non-linear scales, the $\overline{T}_{\textrm{H{i}}}$ degeneracy can potentially be broken [43, 110]. We defer an extended study on this topic involving simulations to future work.
@@
Figure 8 motivates the decision for MeerKLASS to pursue the much larger $10{,}000\,{\rm deg}^{2}$ survey (as is now planned) compared to the proposed $4{,}000\,{\rm deg}^{2}$ in the original white paper [119]. Nevertheless, there are important caveats to adopting a wide-and-thin strategy. First, additive systematics are less effectively averaged down in shallower data, making the survey more vulnerable to residual foregrounds/RFI and instrumental artefacts [91, 145, 68, 48]. Second, wider coverage will inevitably push observations closer to the Galactic plane, where foregrounds are brighter and more complex [135, 72, 79]. This raises the requirement for robust foreground-cleaning methods to perform reliably at lower Galactic latitudes. Both issues will be investigated in detail using ongoing MeerKLASS observations, which will provide an empirical testbed for evaluating the performance of foreground cleaning and systematic control across different regions of sky.
@@
Figure 8: (Forecast improvement on constraints by changing sky area for a fixed observational time (quoted in grey). For $f_{\rm NL}$, > 100\,{>}\,100% gains on the original proposed (nominal [119]) area are available by adopting a “wide and shallow” MeerKLASS survey. Bottom panel shows the absolute error on $f_{\rm NL}$. With Rubin Observatory multi-tracers, Planck18 will be surpassed, and primordial non-Gaussianity detections become viable. Next generation potential is shown in purple shading (following [128]).
@@ 6.2 Probing the growth rate of structure at radio wavelengths: independent tests of gravity
Figure 9: Forecast constraints on growth rate from full MeerKLASS $10{,}000\,{\rm deg}^{2}$ survey, showing its potential to distinguish general relativity (GR) from modified theories of gravity; parameterised by $\mu_{0}$. Constraints are compared to the current state-of-the-art from eBOSS and DESI [66, 63].
Redshift-space distortions (RSD) in the large-scale distribution of matter provide a direct probe into the growth rate of structure, $f(z)$ [77]. Measuring $f(z)$ across redshift enables stringent tests of gravity, since any departure from the predictions of General Relativity would manifest as a change in the growth history. In this way, RSD measurements complement distance-scale probes such as baryon acoustic oscillations (BAO), offering a route to test whether extensions to Einstein’s theory are required [84].
Unlike BAO, which are highly robust to observational systematics [103], RSD measurements are more vulnerable to survey-specific effects [116] that can introduce anisotropies in the clustering signal. This increases the importance of obtaining independent RSD measurements across different experiments, where uncorrelated systematics can be identified and mitigated in combined analyses. Intensity mapping offers such an opportunity: by probing the same underlying large-scale modes with a different set of contaminants and instrumental systematics, MeerKLASS can provide an independent validation of growth rate measurements.
Figure 9 illustrates the forecast performance of MeerKLASS in constraining the growth rate. The plot shows how intensity mapping can act as a powerful discriminator between modified gravity models, here parameterised by $\mu_{0}$, which governs deviations from GR in the effective Newtonian coupling [127, 60]. MeerKLASS is forecast to deliver competitive constraints, exceeding current state-of-the-art spectroscopic surveys such as eBOSS [66] and DESI [63], and extending measurements to complementary redshift ranges. This demonstrates the potential of MeerKLASS as an effective “ruler” on modified gravity.
@@
The main challenge, however, will be to demonstrate sufficient control of observational systematics. Foreground cleaning, beam chromaticity, and survey masks all introduce anisotropies that can mimic or bias an RSD signal if not modelled with high fidelity [54]. Developing robust pipelines that accurately capture these effects will therefore be critical. Future work with MeerKLASS will focus on meeting this challenge, with the goal of contributing a new, independent constraint on the growth rate of structure that will strengthen the global programme of testing gravity with large-scale structure.
@@ 6.3 Multi-tracer cosmology with MeerKLASS for primordial non-Gaussianity (PNG)
Perhaps the most powerful opportunity offered by intensity mapping is its spectroscopic access to ultra-large scales, where few other probes can compete [9]. On these scales, the 21cm line provides a unique window into the physics of the early Universe, particularly primordial non-Gaussianity (PNG) [87, 58]. PNG refers to deviations from Gaussian initial conditions seeded during inflation, which imprint themselves as a scale-dependent modulation of galaxy and Hi bias on the largest modes and is modulated by the parameter $f_{\rm NL}$. Measuring this so-called local-$f_{\rm NL}$ with high precision is one of the most linear methods, thus less complicated by small-scale astrophysics, for distinguishing between competing models of inflation.
@@
The combination of MeerKLASS and early-Rubin is therefore forecast to provide competitive $\sigma(f_{\rm NL})$ values, potentially surpassing the current state-of-the-art CMB constraints from Planck [109]. Looking ahead, the synergy between the full SKAO survey and Rubin could push uncertainties below the crucial threshold of $\sigma(f_{\rm NL})<1$, a milestone that would allow many multi-field inflationary models to be definitively ruled out. Achieving this level of precision would represent a landmark result in fundamental cosmology.
Realising this potential will require stringent control of observational systematics. Foreground cleaning and the associated problem of signal loss are especially critical, since they predominantly affect the large-scale modes that drive PNG sensitivity. However, as shown in Cunnington et al. [57], robust reconstruction of the transfer function offers a viable pathway even for precision measurements of this kind. The incoming MeerKLASS data therefore provide an ideal test-bed for developing and validating these methods, paving the way towards next-generation multi-tracer PNG cosmology.
Additional challenges for local $f_{\rm NL}$ constraints, more general to large-scale structure probes, is the degeneracy with the primordial bias parameter $b_{\phi}$. Detections of $f_{\text{NL}}\,{\neq}\,0$ are thus reliant on a tight prior in the quantity $b_{\phi}f_{\text{NL}}$. As with galaxy surveys, the value of $b_{\phi}$ is not currently well constrained for the Hi tracer [21, 20]. Future work using the latest state-of-the-art hydrodynamical simulations [e.g. 121] can explore this issue further. Our multi-tracer approach will bring benefits to this problem too by reducing the sensitivity to each tracer’s $b_{\phi}$ prior uncertainty [22], especially when tracers differ in redshift distribution and bias evolution, exactly the case for MeerKLASS and Rubin.
@@ 7 Summary
MeerKLASS has established itself as a pioneering programme for validating and exploiting single-dish 21cm intensity mapping with the MeerKAT array. Through careful development of calibration, map-making, and analysis methods, the collaboration has overcome key challenges including foreground contamination, signal loss, and the transformation of survey maps into Cartesian volumes for Fourier analysis. These efforts have culminated in the first robust detections of cosmological clustering in cross-correlation with optical galaxy surveys, alongside significant progress towards recovering auto-correlation measurements and stacked Hi emission.
@@ Appendix A Fisher forecast formalism
so that the $\sigma_{8}$ parameter makes the degeneracy between parameters and power spectrum amplitude explicit. In Equation 19, $\overline{T}_{\textrm{H{i}}}$ is the background mean Hi brightness temperature, $b_{\textrm{H{i}}}$ is the linear bias for Hi, and $f$ is the growth rate of structure. For the Hi bias, we fit to hydrodynamical simulations [144] providing the model
@@
where we use the model for $\Omega_{\textrm{H{i}}}$ based on the one used in SKA Cosmology SWG et al. [128] but adapted based on the latest $\Omega_{\textrm{H{i}}}$ constraints at higher redshift from MeerKLASS intensity mapping [56]
@@
In Equation 19, we include the possibility for local primordial non-Gaussianity (PNG) to influence the power, which occurs by the bias acquiring a scale-dependent correction given by the $b_{\phi}f_{\text{NL}}\mathcal{M}^{-1}$ term. $f_{\text{NL}}$ modulates the PNG contribution and we assume $f_{\text{NL}}\,{=}\,0$ as fiducial default. $b^{\textrm{H{i}}}_{\phi}$ is the PNG bias parameter for Hi, where is common to adopt the standard universality assumption of the halo mass function (which indeed we use for the galaxy power spectrum). However, to be more realistic, we instead interpolate the PNG bias for Hi from the simulation results in Barreira [21], using the model
@@
21cm intensity mapping pixel positions in redshift and angular sky coordinates are measured directly, but converting them into comoving distances requires assuming a fiducial cosmology. If this assumed model differs from the true one, the inferred radial and transverse scales are distorted differently, introducing an artificial anisotropy in the power spectrum. This purely geometric distortion is known as the Alcock–Paczynski (AP) effect [6]. The effect can be described by the rescaling factors present in Equation 19
@@
To correct for any change in volume caused by the different cosmologies, the power spectrum model must be multiplied by $\alpha_{\parallel}^{-1}\alpha_{\perp}^{-2}$, which explains this factor at the front of Equation 19. Lastly, $P_{\text{m}}$ in Equation 19, is the matter power spectrum obtained from a Boltzmann solver [we used CAMB 90, as standard].
@@
where we assume $b_{\text{g}}\,{=}\,1\,{+}\,0.84z$, for a Rubin-like bias [4], and $b_{\text{g}}\,{=}\,2.0,1.2,2.1$ for the DESI LRG, ELG and QSO samples respectively. We adopt the standard universality assumption for the galaxy PNG bias
@@
where $R_{\text{b}}(z)\,{=}\,\chi(z)\,\theta_{\text{FWHM}}(z)/2\sqrt{2\ln 2}$. $\theta_{\text{FWHM}}$ is radio telescope specific but for a generic dish size with diameter $D_{\text{dish}}$ and observations at a frequency $\nu$, we can say $\theta_{\text{FWHM}}\,{=}\,c\,/\,(\nu D_{\text{dish}})$. Within a Fisher forecast or simulation, this beam term is known exactly, and if desired, can be perfectly deconvolved from the power spectrum (as we explore in Appendix B). In real analyses, however, the beam can be more complex and unknown; hence, it would be impractical to deconvolve fully from the data. We therefore leave the beam in the signal and model it to more closely match a real data analysis approach [as discussed in 28].
The noise power in Equation 32 is given just by the telescope’s thermal noise, i.e. we neglect shot-noise, which will be minimal for a Hi intensity mapping survey [144, 131], hence this is modelled as
@@
which follows Cunnington [52] which tuned this model to match recent intensity mapping observations with the MeerKAT pilot survey [145]. The time per pixel $t_{\text{pix}}$ in the noise RMS (Equation 36), is given by the total observation time $t_{\text{obs}}$ from the survey, shared among each pixel. Thus we define it as
@@
in which we are inherently assuming there is no $k$-mode coupling i.e. a Gaussian covariance approximation. For future 21cm intensity mapping surveys this should become a more robust assumption, but for current results this appears not to be true [100, 47]. We leave an exploration into the consequences of wide-spread mode coupling for future work. The number of unique modes is given by
@@
Multipole expansion: It is convenient to adopt the multipole expansion of the anisotropic power spectrum in Equation 19 [54, 28, 130, 29], which will be more representative of the estimators applied to observed data. This expansion is achieved using
@@
Note, in other literature, e.g. Bernal et al. [28], which we otherwise follow closely, the AP pre-factor $\alpha^{-1}_{\parallel}\alpha^{-2}_{\perp}$ is taken outside the integral of Equation 46, but these approaches are identical since the AP parameters are scale-independent and can be taken outside as constants at fixed $z$.
Fisher forecasting: To include our observable in a Fisher forecast, we consider that the set of fluctuations $\delta(\bm{x})$ within a survey map’s pixels with position $\bm{x}$ can be reasonably modelled as a Gaussian distribution with mean of zero and pixel covariance $\textbf{{C}}\,{\equiv}\,\langle\delta^{*}(\bm{x})\delta(\bm{x})\rangle$. The Fisher information matrix is then given by [141, 125, 126]
@@
For the multipole expansion formalism with an observable like $\delta(\bm{x})$, the Fisher matrix simplifies to [140],
@@
which is a matrix of sub-covariance matrices for each multipole combination. Whilst we do not include $k$-mode coupling in these sub-matrices, modelling the covariance between multipoles is necessary for an unbiased analysis [73]. We highlight here, how our approach is avoiding any of the damping terms (e.g. due to the beam) entering the Fisher derivatives $\partial P_{\ell}/\partial\theta_{i}$, which are sensitive only to the shape of the cosmological signal coming from Equation 46 and Equation 19. The loss of information from any instrumental or observational damping will be contained in the covariance (51), through the factor $P_{\text{obs}}$ (Equation 32).
For computing the derivatives for the Fisher matrix, we follow Blanchard et al. [33] and say that given Equation 46 we have
@@
For further reading and related works on forecasting in the context of Hi intensity mapping, we recommend Camera et al. [38], Bull et al. [36], Fonseca et al. [70], Alonso and Ferreira [7], Alonso et al. [9], Pourtsidou et al. [112], Obuljen et al. [106], Bernal et al. [28], Berti et al. [30].

### v2

@@ 6.1.1 Optimising survey strategy by increasing sky area
In this simple test we let the parameters $\bm{\theta}\,{=}\,\{b_{\textrm{H{i}}},f,\alpha_{\perp},\alpha_{\parallel},f_{\text{NL}}\}$ vary, whilst assuming everything else fixed. We extract $H(z)$ from the AP parameter $\alpha_{\parallel}(z)$ (see Equation 27), and marginalise over $b_{\textrm{H{i}}}$ and $\alpha_{\perp}$. The parameters $b_{\textrm{H{i}}}$ and $f$ should be fully degenerate in linear theory for the monopole of the Hi power spectrum. For intensity mapping, it has been shown that going to higher order multipoles is especially crucial for breaking such degeneracies [Bernal:2019jdo, Kennedy:2021srz, Rubiola:2021afc], so in this analysis we include the quadrupole and hexadecapole, i.e. $\bm{P}_{\ell}\,{=}\,\{P_{0},P_{2},P_{4}\}$, which assumes observational anisotropies can be well controlled [Blake:2019ddd, Cunnington:2020mnn, Soares:2020zaq]. There will be additional degeneracies with $\overline{T}_{\textrm{H{i}}}$ and $\sigma_{8}$, for which we fix here, thus we are inherently assuming they have been independently well measured, or alternatively constraints would include these factors as degenerate quantities, e.g. $f\sigma_{8}\overline{T}_{\textrm{H{i}}}$. The $\sigma_{8}$ degeneracy is a general problem for large-scale structure, but the $\overline{T}_{\textrm{H{i}}}$ is, of course, more unique to intensity mapping [identified in Bernal:2019jdo, Castorina:2019zho]. Exploiting tomography, i.e. leveraging the differing redshift evolution of degenerate parameters across bins, can provide some immediate relief here. Additionally, 1-point statistics, such as the voxel intensity distribution (VID) [Breysse:2016szq, Sato-Polito:2022fkd, Bernal:2023ovz], can capture the non-Gaussian information in the line intensity maps, thus combining the VID with the power spectrum helps to break degeneracies between astrophysical and cosmological parameters [Sabla:2024dxz]. Furthermore, going to non-linear scales with interferometers, where scale dependence will enter and since Hi clustering should be accurately described by perturbative methods at these mildly non-linear scales, the $\overline{T}_{\textrm{H{i}}}$ degeneracy can potentially be broken [Castorina:2019zho, Pourtsidou:2022gsb]. We defer an extended study on this topic involving simulations to future work.
@@
Figure 9 motivates the decision for MeerKLASS to pursue the much larger $10{,}000\,{\rm deg}^{2}$ survey (as is now planned) compared to the proposed $4{,}000\,{\rm deg}^{2}$ in the original white paper [MeerKLASS:2017vgf]. Nevertheless, there are important caveats to adopting a wide-and-thin strategy. First, additive systematics are less effectively averaged down in shallower data, making the survey more vulnerable to residual foregrounds/RFI and instrumental artefacts [Li:2020bcr, Wang:2020lkn, Engelbrecht:2024eoc, Chen:2025bcx]. Second, wider coverage will inevitably push observations closer to the Galactic plane, where foregrounds are brighter and more complex [Staveley-Smith:2016gst, 2024MNRAS.531..649G, Irfan:2021xuk]. This raises the requirement for robust foreground-cleaning methods to perform reliably at lower Galactic latitudes. Both issues will be investigated in detail using ongoing MeerKLASS observations, which will provide an empirical testbed for evaluating the performance of foreground cleaning and systematic control across different regions of sky.
@@
Figure 9: Forecast improvement on constraints by changing sky area for a fixed observational time (quoted in grey). For $f_{\rm NL}$, > 100\,{>}\,100% gains on the original proposed [nominal MeerKLASS:2017vgf] area are available by adopting a “wide and shallow” MeerKLASS survey. Bottom panel shows the absolute error on $f_{\rm NL}$. A multi-tracer approach with a Rubin Observatory galaxy clustering probe (green-dashed line) will see Planck18 surpassed, and primordial non-Gaussianity detections become viable. Next generation potential is shown in purple shading [following SKA:2018ckk]. We assume the full Y10 gold sample for Rubin with the SKAO, but a restricted version for MeerKLASS (details given in the text).
@@ 6.2 Probing the growth rate of structure at radio wavelengths: independent tests of gravity
Figure 10: Forecast constraints on growth rate from full MeerKLASS $10{,}000\,{\rm deg}^{2}$ survey, showing its potential to distinguish general relativity (GR) from modified theories of gravity; parameterised by $\mu_{0}$. Constraints are compared to the current state-of-the-art from eBOSS and DESI [eBOSS:2020yzd, DESI:2024jxi].
Redshift-space distortions (RSD) in the large-scale distribution of matter provide a direct probe into the growth rate of structure, $f(z)$ [Huterer:2022dds]. Measuring $f(z)$ across redshift enables stringent tests of gravity, since any departure from the predictions of General Relativity would manifest as a change in the growth history. In this way, RSD measurements complement distance-scale probes such as baryon acoustic oscillations (BAO), offering a route to test whether extensions to Einstein’s theory are required [Joyce:2014kja].
Unlike BAO, which are highly robust to observational systematics [eBOSS:2021rwq], RSD measurements are more vulnerable to survey-specific effects [Samushia:2011cs] that can introduce anisotropies in the clustering signal. This increases the importance of obtaining independent RSD measurements across different experiments, where uncorrelated systematics can be identified and mitigated in combined analyses. Intensity mapping offers such an opportunity: by probing the same underlying large-scale modes with a different set of contaminants and instrumental systematics, MeerKLASS can provide an independent validation of growth rate measurements.
Figure 10 illustrates the forecast performance of MeerKLASS in constraining the growth rate. The plot shows how intensity mapping can act as a powerful discriminator between modified gravity models, here parameterised by $\mu_{0}$, which governs deviations from GR in the effective Newtonian coupling [Simpson:2012ra, DES:2018ufa]. MeerKLASS is forecast to deliver competitive constraints, exceeding current state-of-the-art spectroscopic surveys such as eBOSS [eBOSS:2020yzd] and DESI [DESI:2024jxi], and extending measurements to complementary redshift ranges. This demonstrates the potential of MeerKLASS as an effective “ruler” on modified gravity.
@@
The main challenge, however, will be to demonstrate sufficient control of observational systematics. Foreground cleaning, beam chromaticity, and survey masks all introduce anisotropies that can mimic or bias an RSD signal if not modelled with high fidelity [Cunnington:2020mnn]. Developing robust pipelines that accurately capture these effects will therefore be critical. Future work with MeerKLASS will focus on meeting this challenge, with the goal of contributing a new, independent constraint on the growth rate of structure that will strengthen the global programme of testing gravity with large-scale structure.
@@ 6.3 Multi-tracer cosmology with MeerKLASS for primordial non-Gaussianity (PNG)
Perhaps the most powerful opportunity offered by intensity mapping is its spectroscopic access to ultra-large scales, where few other probes can compete [Alonso:2015uua]. On these scales, the 21cm line provides a unique window into the physics of the early Universe, particularly primordial non-Gaussianity (PNG) [Komatsu:2001rj, Dalal:2007cu]. PNG refers to deviations from Gaussian initial conditions seeded during inflation, which imprint themselves as a scale-dependent modulation of galaxy and Hi bias on the largest modes and is modulated by the parameter $f_{\rm NL}$. Measuring this so-called local-$f_{\rm NL}$ with high precision is one of the most linear methods, thus less complicated by small-scale astrophysics, for distinguishing between competing models of inflation.
@@
The combination of MeerKLASS and early-Rubin is therefore forecast to provide competitive $\sigma(f_{\rm NL})$ values, potentially surpassing the current state-of-the-art CMB constraints from Planck [Planck:2018vyg]. Looking ahead, the synergy between the full SKAO survey and Rubin could push uncertainties below the crucial threshold of $\sigma(f_{\rm NL})<1$, a milestone that would allow many multi-field inflationary models to be definitively ruled out. Achieving this level of precision would represent a landmark result in fundamental cosmology.
Realising this potential will require stringent control of observational systematics. Foreground cleaning and the associated problem of signal loss are especially critical, since they predominantly affect the large-scale modes that drive PNG sensitivity. However, as shown in Cunnington:2023jpq, robust reconstruction of the transfer function offers a viable pathway even for precision measurements of this kind. The incoming MeerKLASS data therefore provide an ideal test-bed for developing and validating these methods, paving the way towards next-generation multi-tracer PNG cosmology.
Additional challenges for local $f_{\rm NL}$ constraints, more general to large-scale structure probes, is the degeneracy with the primordial bias parameter $b_{\phi}$. Detections of $f_{\text{NL}}\,{\neq}\,0$ are thus reliant on a tight prior in the quantity $b_{\phi}f_{\text{NL}}$. As with galaxy surveys, the value of $b_{\phi}$ is not currently well constrained for the Hi tracer [Barreira:2021dpt, Barreira:2022sey]. Future work using the latest state-of-the-art hydrodynamical simulations [e.g. Schaye:2025xuv] can explore this issue further. Our multi-tracer approach will bring benefits to this problem too by reducing the sensitivity to each tracer’s $b_{\phi}$ prior uncertainty [Barreira:2023rxn], especially when tracers differ in redshift distribution and bias evolution, exactly the case for MeerKLASS and Rubin.
@@ 7 Summary
MeerKLASS has established itself as a pioneering programme for validating and exploiting single-dish 21cm intensity mapping with the MeerKAT array. Through careful development of calibration, map-making, and analysis methods, the collaboration has overcome key challenges including foreground contamination, signal loss, and the transformation of survey maps into Cartesian volumes for Fourier analysis. These efforts have culminated in the first robust detections of cosmological clustering in cross-correlation with optical galaxy surveys, alongside significant progress towards recovering auto-correlation power spectra measurements and stacked Hi emission.
@@ Appendix A Fisher forecast formalism
so that the $\sigma_{8}$ parameter makes the degeneracy between parameters and power spectrum amplitude explicit. In Equation 20, $\overline{T}_{\textrm{H{i}}}$ is the background mean Hi brightness temperature, $b_{\textrm{H{i}}}$ is the linear bias for Hi, and $f$ is the growth rate of structure. For the Hi bias, we fit to hydrodynamical simulations [Villaescusa-Navarro:2018vsg] providing the model
@@
where we use the model for $\Omega_{\textrm{H{i}}}$ based on the one used in SKA:2018ckk but adapted based on the latest $\Omega_{\textrm{H{i}}}$ constraints at higher redshift from MeerKLASS intensity mapping [Cunnington:2022uzo]
@@
In Equation 20, we include the possibility for local primordial non-Gaussianity (PNG) to influence the power, which occurs by the bias acquiring a scale-dependent correction given by the $b_{\phi}f_{\text{NL}}\mathcal{M}^{-1}$ term. $f_{\text{NL}}$ modulates the PNG contribution and we assume $f_{\text{NL}}\,{=}\,0$ as fiducial default. $b^{\textrm{H{i}}}_{\phi}$ is the PNG bias parameter for Hi, where is common to adopt the standard universality assumption of the halo mass function (which indeed we use for the galaxy power spectrum). However, to be more realistic, we instead interpolate the PNG bias for Hi from the simulation results in Barreira:2021dpt, using the model
@@
21cm intensity mapping pixel positions in redshift and angular sky coordinates are measured directly, but converting them into comoving distances requires assuming a fiducial cosmology. If this assumed model differs from the true one, the inferred radial and transverse scales are distorted differently, introducing an artificial anisotropy in the power spectrum. This purely geometric distortion is known as the Alcock–Paczynski (AP) effect [Alcock:1979mp]. The effect can be described by the rescaling factors present in Equation 20
@@
To correct for any change in volume caused by the different cosmologies, the power spectrum model must be multiplied by $\alpha_{\parallel}^{-1}\alpha_{\perp}^{-2}$, which explains this factor at the front of Equation 20. Lastly, $P_{\text{m}}$ in Equation 20, is the matter power spectrum obtained from a Boltzmann solver [we used CAMB Lewis:1999bs, as standard].
@@
where we assume $b_{\text{g}}\,{=}\,1\,{+}\,0.84z$, for a Rubin-like bias [LSSTScience:2009jmu], and $b_{\text{g}}\,{=}\,2.0,1.2,2.1$ for the DESI LRG, ELG and QSO samples respectively. We adopt the standard universality assumption for the galaxy PNG bias
@@
where $R_{\text{b}}(z)\,{=}\,\chi(z)\,\theta_{\text{FWHM}}(z)/2\sqrt{2\ln 2}$. $\theta_{\text{FWHM}}$ is radio telescope specific but for a generic dish size with diameter $D_{\text{dish}}$ and observations at a frequency $\nu$, we can say $\theta_{\text{FWHM}}\,{=}\,c\,/\,(\nu D_{\text{dish}})$. Within a Fisher forecast or simulation, this beam term is known exactly, and if desired, can be perfectly deconvolved from the power spectrum (as we explore in Appendix B). In real analyses, however, the beam can be more complex and unknown; hence, it would be impractical to deconvolve fully from the data. We therefore leave the beam in the signal and model it to more closely match a real data analysis approach [as discussed in Bernal:2019jdo].
The noise power in Equation 33 is given just by the telescope’s thermal noise, i.e. we neglect shot-noise, which will be minimal for a Hi intensity mapping survey [Villaescusa-Navarro:2018vsg, Spinelli:2019smg], hence this is modelled as
@@
which follows Cunnington:2022ryj which tuned this model to match recent intensity mapping observations with the MeerKAT pilot survey [Wang:2020lkn]. The time per pixel $t_{\text{pix}}$ in the noise RMS (Equation 37), is given by the total observation time $t_{\text{obs}}$ from the survey, shared among each pixel. Thus we define it as
@@
in which we are inherently assuming there is no $k$-mode coupling i.e. a Gaussian covariance approximation. For future 21cm intensity mapping surveys this should become a more robust assumption, but for current results this appears not to be true [MeerKLASS:2024ypg, Chen:2025til]. We leave an exploration into the consequences of wide-spread mode coupling for future work. The number of unique modes is given by
@@
Multipole expansion: It is convenient to adopt the multipole expansion of the anisotropic power spectrum in Equation 20 [Cunnington:2020mnn, Bernal:2019jdo, Soares:2020zaq, Berti:2022ilk], which will be more representative of the estimators applied to observed data. This expansion is achieved using
@@
Note, in other literature, e.g. Bernal:2019jdo, which we otherwise follow closely, the AP pre-factor $\alpha^{-1}_{\parallel}\alpha^{-2}_{\perp}$ is taken outside the integral of Equation 47, but these approaches are identical since the AP parameters are scale-independent and can be taken outside as constants at fixed $z$.
Fisher forecasting: To include our observable in a Fisher forecast, we consider that the set of fluctuations $\delta(\bm{x})$ within a survey map’s pixels with position $\bm{x}$ can be reasonably modelled as a Gaussian distribution with mean of zero and pixel covariance $\textbf{{C}}\,{\equiv}\,\langle\delta^{*}(\bm{x})\delta(\bm{x})\rangle$. The Fisher information matrix is then given by [Tegmark:1997rp, Seo:2003pu, Seo:2007ns]
@@
For the multipole expansion formalism with an observable like $\delta(\bm{x})$, the Fisher matrix simplifies to [Taruya:2011tz],
@@
which is a matrix of sub-covariance matrices for each multipole combination. Whilst we do not include $k$-mode coupling in these sub-matrices, modelling the covariance between multipoles is necessary for an unbiased analysis [Grieb:2015bia]. We highlight here, how our approach is avoiding any of the damping terms (e.g. due to the beam) entering the Fisher derivatives $\partial P_{\ell}/\partial\theta_{i}$, which are sensitive only to the shape of the cosmological signal coming from Equation 47 and Equation 20. The loss of information from any instrumental or observational damping will be contained in the covariance (52), through the factor $P_{\text{obs}}$ (Equation 33).
For computing the derivatives for the Fisher matrix, we follow Euclid:2019clj and say that given Equation 47 we have
@@
For further reading and related works on forecasting in the context of Hi intensity mapping, we recommend Camera:2013kpa, Bull:2014rha, Fonseca:2015laa, Alonso:2015sfa, Alonso:2015uua, Pourtsidou:2016dzn, Obuljen:2017jiy, Bernal:2019jdo, Berti:2023viz.